10-3 Direct Labor Variances [L03]
SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company's products is
grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week,
the company prepared 4,000 of these meals using 960 direct labor-hours. The company paid these direct labor workers a total of $9,600 for this work, or $10.00 per hour. According to the standard cost card for this meal, it should require 0.25 direct labor-hours at a cost of $9.75 per hour. -

Required:
l. What direct labor cost should have been incurred to prepare 4,000 meals? How much does this differ from the actual direct labor cost?

2. Break down the difference computed in (1) above into a labor rate variance and a labor efficiency variance.
Answer #1
Standard timeto prepare a meal =0.25 labor hours
Total meals prepared=4000
Standard time required to prepare 4000 meals=
4000×0.25=1000 labor hours
Standard labor hours required =1000 hrs
Standard labor hour rate=9.75 per hour
Standard cost to prepare 4000 meals
Standard hrs×Standard labor hour rate
1000×9.75=9750
Difference in cost
Standard cost-Actual cost
9750-9600=150
Labor rate variance
Actual labor hours×(Standard rate-Actual rate)
Actual labor hours=960
Standard rate =9.75 per hr
Actual per =10 per hr
960×(9.75-10)=240(unfavorable)
Labor efficiency variance
Standard hourly rate×(standard hours-Actual hours)
9.75×(1000-960)=390(favourable)
Labor rate variance        =240(unfavourable)
Labor efficiency variance =390(favourable)
Difference                       =150(favourable)